Two point charges are fixed on the y axis: a negative point charge q1 = -26 μC at y1 = +0.23 m and a positive point charge q2 at y2 = +0.33 m. A third point charge q = +8.7 μC is fixed at the origin. The net electrostatic force exerted on the charge q by the other two charges has a magnitude of 28 N and points in the +y direction. Determine the magnitude of q2.
Answer #1

q2 F1-F2-28N 0.33 q1- 26uC F2 F1 0.23m q-8.7uC

F_{1}=kqq_{1}/r^{2}

Fi9 10 8.7 102.610C/(0.23m) 2

F1-9 108.7 10 2.6100.23

F1- 38.48N

==============

F_{2}=kqq_{2}/r^{2}

F2 9 10 8.7 10 10/(0.33m)

F2 9 10 8.7 10100.33

F, 0.719 *

=================

Given Net Force=28N

so F1-F2 = 28N

38.48N-0.719*q_{2}=28N

38.48-28=0.719*q_{2}

(38.48-28)/0.719=q_{2}

ANSWER: {color{Red} q_{2}=14.58mu C}

==========================